Suppose that the management of Sun City resort is trying to understand the income elasticity of their resort. As an advisor, they provide you with the following figures;

If the average income of visitors is R3 000 (which is the current income), then 45 000 visitors visit Sun City resort each year and if the average income increases to R5 000, then 60 000 visitors visit Sun City resort each year.


  1. Calculate the income elasticity of demand using the arc method.
  2. Based on your answer above (1), explain what type of a product (service) the resort is.
1
Expert's answer
2021-09-06T07:08:19-0400

1)

Arc method:

Income elasticity of demand"\frac{{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}}{{\frac{I_2-I_1}{\frac{I_2+I_1}{2}}}}"

"Q_1=45,000\\Q_2=60,000\\I_1=R3,000\\I_2=R5,000"

Income elasticity of demand "\frac{{\frac{60000-45000}{\frac{60000+45000}{2}}}}{{\frac{5000-3000}{\frac{5000+3000}{2}}}}"

"=\frac{{\frac{15000}{\frac{105000}{2}}}}{{\frac{2000}{\frac{8000}{2}}}}"

"=\frac{{\frac{15000}{52500}}}{{\frac{2000}{4000}}}\\=\frac{15000}{52500}\times \frac{4000}{2000}\\=0.57"


Therefore, the income elasticity of demand is 0.57.


2)

Normal goods: - if the quantity demanded of any good increases with the increase in consumer’s income, these types of goods are known as normal goods. The income elasticity of these goods is positive because these goods have a direct relationship with income which means the quantity demanded of normal goods increases with the increase in income and decreases with the decrease in income.

Since the income elasticity of the resort is positive so the resort is normal good.